# AGC-Drive: A Large-Scale Dataset for Real-World Aerial-Ground Collaboration in Driving Scenarios

### v1

@@ Abstract
By sharing information across multiple agents, collaborative perception helps autonomous vehicles mitigate occlusions and improve overall perception accuracy. While most previous work focus on vehicle-to-vehicle and vehicle-to-infrastructure collaboration, with limited attention to aerial perspectives provided by UAVs, which uniquely offer dynamic, top-down views to alleviate occlusions and monitor large-scale interactive environments. A major reason for this is the lack of high-quality datasets for aerial-ground collaborative scenarios. To bridge this gap, we present AGC-Drive, the first large-scale real-world dataset for Aerial-Ground Cooperative 3D perception. The data collection platform consists of two vehicles, each equipped with five cameras and one LiDAR sensor, and one UAV carrying a forward-facing camera and a LiDAR sensor, enabling comprehensive multi-view and multi-agent perception. Consisting of approximately 120K LiDAR frames and 440K images, the dataset covers 14 diverse real-world driving scenarios, including urban roundabouts, highway tunnels, and on/off ramps. Notably, 19.5% of the data comprises dynamic interaction events, including vehicle cut-ins, cut-outs, and frequent lane changes. AGC-Drive contains 400 scenes, each with approximately 100 frames and fully annotated 3D bounding boxes covering 13 object categories. We provide benchmarks for two 3D perception tasks: vehicle-to-vehicle collaborative perception and vehicle-to-UAV collaborative perception. Additionally, we release an open-source toolkit, including spatiotemporal alignment verification tools, multi-agent visualization systems, and collaborative annotation utilities. The dataset and code are available at https://github.com/PercepX/AGC-Drive.
@@ 1 Introduction
Unlike V2V and V2I systems, Aerial-Ground Cooperative Perception (AGCP) introduces overhead Unmanned Aerial Vehicle (UAV)-based sensing, offering dynamic, adaptive, and high-altitude global perspectives to complement ground-based systems. UAVs can flexibly cover target areas, dynamically alleviate perception blind spots, enhance long-range target observation, and improve multi-object occlusion reasoning in complex traffic environments. This makes AGCP a valuable complement to existing V2X systems, particularly in scenarios involving open roads, dynamic intersections, dense traffic, and emergency scenarios.
@@
Although several datasets [13, 14, 15, 16, 12] have introduced UAVs into collaborative perception, they either focus on UAV-to-UAV cooperation [13, 14] or are collected in simulated environments [13, 14, 15, 12]. The only available real-world dataset [16] provides 2D annotations, but it is not collected in driving scenarios. Due to hardware limitations, none of these existing datasets include LiDAR-equipped UAVs. To bridge this gap, we introduce AGC-Drive, a dataset collected by a system consisting of two vehicles and one UAV (Figure 1) over approximately three months. From over 80 hours of collected data, we selected 400 sequences covering 14 scene categories. Each sequence contains 100 sets of data sampled at 10 Hz, with each set including 14 LiDAR frames and image frames, resulting in over 1.6M annotated 3D bounding boxes. Notably, we carefully designed the data collection routes to ensure a wide range of road conditions and vehicle interaction scenarios, covering high-risk environments such as urban roundabouts, highway tunnels, on/off ramps, and rural construction zones. Approximately 18.2% of the data involves dynamic events like vehicle cut-ins, cut-outs, and dense lane changes.
@@
  • We provide over 1.6M annotated 3D bounding boxes for 13 categories, covering 120K LiDAR frames and 440K multi-view images, collected across 14 types of road environments and dynamic interaction scenarios such as lane changes and overtaking.
@@ 2 Related work
Table 1: A detailed comparison between representative single-UAV/cooperative perception datasets. Notes: V = Vehicle, I = Infrastructure, U = UAV. "V2V&I" indicates V2V and V2I, "V2V&U" indicates V2V and V2U. "MvCamera" denotes a Multi-view Camera System, "UAV-L" represents a UAV equipped with LiDAR, and "Driving" represents whether the dataset is collected in autonomous driving scenarios.
Mode Dataset Year Source Agent Sensor scenario types 3D boxes Classes MvCams Driving UAV-L
@@
V2V&U AGC-Drive(Ours) 2025 Real Veh & UAV C & L & R 14 1.6M 13
@@ 3.1 Setup
2*Vehicle LiDAR RoboSense Ruby Plus(*1) 128 beams, 10Hz capture frequency, 360°horizontal FOV, -25°to +15°vertical FOV, < 200m range
Camera Sensing Cmaera(*5) front-wide: SG8S-AR0820C-5300-G2A-Hxxx, 8MP, HFOV30°, front-broad: SG8S-AR0820C-5300-G2A-Hxxx, 8MP, HFOV120°, left&right: SG2-AR0231C-0202-GMSL-Hxxx, 2MP, HFOV100°, back: SG2-AR0233C-5200-G2A-Hxxx, 2mp, HFOV121°
@@
UAV LiDAR RoboSense Helios32(*1) 32 beams, 10Hz capture frequency, 360°horizontal FOV, -55°to +15°vertical FOV,< 150m range
@@
Figure 2: Distribution of Driving Environment and Scenario Types.The inner ring categorizes the data into three primary driving environments: Urban (77.5%), Highway (13.0%), and Rural (9.5%). The outer ring further subdivides these environments into specific scenario types.
@@ 3.2 Data Acquisition
Collection. We deployed two equipped vehicles and a UAV operated by professional pilots to collect synchronized data. Each agent locally records data frames, with timestamps aligned to its GPS time, achieving unified time sources across all agents. Data were collected along predefined routes over 2 months under varying day and night conditions, accumulating over 80 hours of driving logs. From the raw data, we curated 400 representative segments (10s each) across 14 scenarios, resulting in 120K LiDAR frames and 440K camera images with annotations. Keyframes were extracted at 10 Hz to construct AGC-V2V, while segments involving UAV participation were used to create AGC-VUC for vehicle-UAV collaborative perception tasks. Additionally, we plan to extracted UAV LiDAR data to support UAV-based 3D perception research.
@@
Figure 3: (a) Point-to-point residual heatmap of the registration results. As indicated by the color bar, colors transition from blue to red, representing registration accuracy from high to low. (b) A zoomed-in view of the dense object region highlighted by the dashed box in (a). (c) Scatter plot of the average residuals for 4000 randomly sampled points, where both the drone and vehicle point clouds are registered to the ego vehicle’s point cloud.
@@ 4 Task And Benchmark
Dataset Format and Split. Our AGC-Drive dataset follows the OPV2V [1] format, containing 14 road types with multiple scenes per type (distribution shown in Fig. 2). Each scene consists of 100 groups of synchronized multi-agent data, collected by 1 drone and 2 vehicles equipped with a total of 14 sensors. In total, we collect 400 scenes, with 320 for training, 90 for validation, and 90 for testing.
@@
  • Upper-bound: Directly shares raw point cloud data among agents before feature extraction.
  • Lower-bound: Independently detects objects and shares detection results among agents.
@@
Experiments Details. For all tasks, the models are adapted to support the AGC-V2V and AGC-VUC data format and sensor configurations. We adopt a unified BEV (Bird’s Eye View) representation. The perception range is set to $[-150\,\text{m},150\,\text{m}]$ along both the X and Y axes. All baseline models were trained with a batch size of 4 for 40 epochs, with a per-GPU batch size of 1 on a computing server equipped with 8 Nvidia L40 GPUs. We use the Adam optimizer with an initial learning rate of 0.001 and apply a cosine learning rate schedule. Each training run takes approximately 6 hours.
Finally, we report standard evaluation metrics following the protocols of nuScenes [18] and OPV2V [1], including mAP@0.5 and mAP@0.7 for 3D object detection, along with $\Delta_{\mathrm{UAV}}$, to quantify the impact of the UAV’s participation on the perception performance of the cooperative system.
@@ 4.1 Benchmark for V2V 3D object detection
Problem Definition and AGC-V2V. The goal of this task is to perform cooperative 3D object detection by leveraging information from multiple connected vehicles to enhance perception performance, especially in challenging scenarios such as long-range detection and occlusions. We construct the AGC-V2V benchmark by selecting collaborative scenes from the AGC-Drive dataset where two vehicles participate without UAV involvement. Each vehicle is equipped with a LiDAR sensor and five cameras. A total of 400 scenes, comprising 80K frames, are included for this benchmark.
Quantitative Results. Table 3 reports the results of V2V 3D object detectio on AGC-V2V. It is unsurprising that the Upper-bound early fusion achieves superior results across all metrics, given its access to fully aggregated information. The overall performance of all frameworks remains relatively low, especially for feature-sensitive intermediate fusion methods, primarily due to the lack of handling for time delays and pose estimation errors. The Lower-bound late fusion performs the worst, due to the accumulation of multiple errors introduced during the collaborative fusion process.
@@
Late Lower-bound 26.6 12.2
Intermediate V2VNet [1] 27.2 14.1
Cobevt [28] 28.6 14.9
Where2comm [13] 34.8 22.5
V2X-ViT [4] 32.7 18.7
Early Upper-bound 58.2 43.1
@@
4.2 Benchmark for VUC3D object cooperation
Problem Definition and AGC-VUC. This task aims to improve 3D object detection by leveraging the complementary perception capabilities of ground vehicles and UAVs. Unlike traditional V2V cooperation, the UAV provides a global top-down view that enhances occlusion handling and long-range detection. To support this, we construct the AGC-VUC benchmark by selecting cooperative scenarios in AGC-Drive that involve two vehicles and one UAV. Each scenario lasts approximately 10 seconds, sampled at 10Hz, resulting in 200 cooperative sequences.
Quantitative Results. Table 4 presents the benchmark results after incorporating the UAV into the cooperative system. Here, we add a new column, denoted as $\Delta_{\mathrm{UAV}}$, to quantify the impact of the UAV’s participation on the perception performance of the cooperative system. It can be observed that the introduction of the UAV improves the performance of all cooperative frameworks across all metrics, except for the Lower-bound. Among them, the Upper-bound achieves the largest improvement, with a $\Delta_{\mathrm{UAV}}$ value of +3.3.
@@ 4.2 Benchmark for VUC3D object cooperation
Co-Mode Model mAP@0.5 mAP@0.7 $\Delta_{\mathrm{UAV}}$
Late Lower-bound 24.9 11.6 -1.0
Intermediate V2VNet [1] 28.5 14.9 +1.0
Cobevt [28] 30.7 16.1 +1.6
Where2comm [13] 36.4 23.1 +1.1
V2X-ViT [4] 33.9 18.4 +0.4
Early Upper-bound 61.3 46.5 +3.3

### v2

@@ Abstract
By sharing information across multiple agents, collaborative perception helps autonomous vehicles mitigate occlusions and improve overall perception accuracy. While most previous work focus on vehicle-to-vehicle and vehicle-to-infrastructure collaboration, with limited attention to aerial perspectives provided by UAVs, which uniquely offer dynamic, top-down views to alleviate occlusions and monitor large-scale interactive environments. A major reason for this is the lack of high-quality datasets for aerial-ground collaborative scenarios. To bridge this gap, we present AGC-Drive, the first large-scale real-world dataset for Aerial-Ground Cooperative 3D perception. The data collection platform consists of two vehicles, each equipped with five cameras and one LiDAR sensor, and one UAV carrying a forward-facing camera and a LiDAR sensor, enabling comprehensive multi-view and multi-agent perception. Consisting of approximately 80K LiDAR frames and 360K images, the dataset covers 14 diverse real-world driving scenarios, including urban roundabouts, highway tunnels, and on/off ramps. Notably, 17% of the data comprises dynamic interaction events, including vehicle cut-ins, cut-outs, and frequent lane changes. AGC-Drive contains 350 scenes, each with approximately 100 frames and fully annotated 3D bounding boxes covering 13 object categories. We provide benchmarks for two 3D perception tasks: vehicle-to-vehicle collaborative perception and vehicle-to-UAV collaborative perception. Additionally, we release an open-source toolkit, including spatiotemporal alignment verification tools, multi-agent visualization systems, and collaborative annotation utilities. The dataset and code are available at https://github.com/PercepX/AGC-Drive.
@@ 1 Introduction
Unlike V2V and V2I systems, Aerial-Ground Cooperative (AGC) Perception introduces overhead Unmanned Aerial Vehicle (UAV)-based sensing, offering dynamic, adaptive, and high-altitude global perspectives to complement ground-based systems. UAVs can flexibly cover target areas, dynamically alleviate perception blind spots, enhance long-range target observation, and improve multi-object occlusion reasoning in complex traffic environments. This makes AGC perception a valuable complement to existing V2X systems, particularly in scenarios involving open roads, dynamic intersections, dense traffic, and emergency scenarios.
@@
Although several datasets [13, 14, 15, 16, 12] have introduced UAVs into collaborative perception, they either focus on UAV-to-UAV cooperation [13, 14] or are collected in simulated environments [13, 14, 15, 12]. The only available real-world dataset [16] provides 2D annotations, but it is not collected in driving scenarios. Due to hardware limitations, none of these existing datasets include LiDAR-equipped UAVs. To bridge this gap, we introduce AGC-Drive, a dataset collected by a system consisting of two vehicles and one UAV (Figure 1) over approximately three months. From over 80 hours of collected data, we selected 350 sequences covering 14 scene categories. Each sequence contains 100 sets of data sampled at 10 Hz, with each set including 14 LiDAR frames and image frames, resulting in over 720K annotated 3D bounding boxes. Notably, we carefully designed the data collection routes to ensure a wide range of road conditions and vehicle interaction scenarios, covering high-risk environments such as urban roundabouts, highway tunnels, on/off ramps, and rural construction zones. Approximately 17% of the data involves dynamic events like vehicle cut-ins, cut-outs, and dense lane changes.
@@
  • We provide over 720K annotated 3D bounding boxes for 13 categories, covering 80K LiDAR frames and 360K multi-view images, collected across 14 types of road environments and dynamic interaction scenarios such as lane changes and overtaking.
@@ 2 Related work
Table 1: Comparison of representative single-UAV and cooperative perception datasets. V/Veh = Vehicle, I/Inf = Infrastructure, U = UAV. "V2V&I" = V2V+V2I, "V2V&U" = V2V+V2U. C, L, R = Camera, Lidar, Radar. "MvCamera" = Multi-view Camera, "UAV-L" = UAV with LiDAR.
Mode Dataset Year Source Agent Sensor
scenario
types
3D boxes Classes MvCams
Driving
UAV-L
@@
V2V&U AGC-Drive(Ours) 2025 Real Veh & UAV C & L & R 14 720K 13
@@ 3.1 Setup
2*Vehicle LiDAR RoboSense Ruby Plus(*1)
128 beams, 10Hz capture frequency, 360°horizontal FOV,
-25°to +15°vertical FOV, < 200m range
Camera Sensing Cmaera(*5)
front-wide: SG8S-AR0820C-5300-G2A-Hxxx, 8MP, HFOV30°,
front-broad: SG8S-AR0820C-5300-G2A-Hxxx, 8MP, HFOV120°,
left&right: SG2-AR0231C-0202-GMSL-Hxxx, 2MP, HFOV100°,
back: SG2-AR0233C-5200-G2A-Hxxx, 2mp, HFOV121°
@@
UAV LiDAR RoboSense Helios32(*1)
32 beams, 10Hz capture frequency, 360°horizontal FOV,
-55°to +15°vertical FOV,< 150m range
@@
Figure 2: Distribution of Driving Environment and Scenario Types.
@@ 3.2 Data Acquisition
Collection. We deployed two equipped vehicles and a UAV operated by professional pilots to collect synchronized data. Each agent locally records data frames, with timestamps aligned to its GPS time, achieving unified time sources across all agents. Data were collected along predefined routes over 2 months under varying day and night conditions, accumulating over 80 hours of driving logs. From the raw data, we curated 350 representative segments (10s each) across 14 scenarios, resulting in 80K LiDAR frames and 360K camera images with annotations. Keyframes were extracted at 10 Hz to construct AGC-V2V, while segments involving UAV participation were used to create AGC-VUC for vehicle-UAV collaborative perception tasks. Additionally, we plan to extracted UAV LiDAR data to support UAV-based 3D perception research.
@@
Figure 3: (a) Residual heatmap of registration results (blue: high accuracy, red: low accuracy). (b) Zoomed view of the dense object region in (a). (c) Scatter plot of average residuals for 4000 randomly sampled points after registering both drone and vehicle point clouds to the ego vehicle.
@@ 4 Task and Benchmark
Dataset Format and Split. Our AGC-Drive dataset follows the OPV2V [1] format, containing 14 road types with multiple scenes per type (distribution shown in Fig. 2). Each scene consists of 100 groups of synchronized multi-agent data, collected by 1 drone and 2 vehicles equipped with a total of 14 sensors. In total, we collect 350 scenes, with 320 for training, 30 for validation and testing.
@@
  • Late Fusion: Directly shares raw point cloud data among agents before feature extraction.
  • Early Fusion: Independently detects objects and shares detection results among agents.
@@
Experiments Details. For all tasks, the models are adapted to support the AGC-V2V and AGC-VUC data format and sensor configurations. We adopt a unified BEV (Bird’s Eye View) representation. The perception range is set to $140.8\,\text{m},140.8\,\text{m}]$ along both the X and Y axes. All baseline models were trained with a batch size of 4 for 60 epochs, with a per-GPU batch size of 1 on a computing server equipped with 8 Nvidia L40 GPUs. We use the Adam optimizer with an initial learning rate of 0.001 and apply a cosine learning rate schedule. Each training run takes approximately 6 hours.
Following the evaluation protocols of nuScenes [18] and OPV2V [1], we report standard metrics including mAP@0.5 and mAP@0.7 for 3D object detection. Furthermore, to analyze the influence of UAV participation, we define the metric $\Delta_{\mathrm{UAV}}$ as:
$\Delta_{\mathrm{UAV}}=\frac{1}{2}\left[\left(m_{0.5}^{V2U}-m_{0.5}^{V2V}\right)+\left(m_{0.7}^{V2U}-m_{0.7}^{V2V}\right)\right],$
which represents the average performance improvement achieved by incorporating aerial perception.
@@ 4.1 Benchmark for V2V 3D object detection
Problem Definition and AGC-V2V. The goal of this task is to perform cooperative 3D object detection by leveraging information from multiple connected vehicles to enhance perception performance, especially in challenging scenarios such as long-range detection and occlusions. We construct the AGC-V2V benchmark by selecting collaborative scenes from the AGC-Drive dataset where two vehicles participate without UAV involvement. Each vehicle is equipped with a LiDAR sensor and five cameras. A total of 350 scenes, comprising 70K frames, are included for this benchmark.
Quantitative Results. Table 3 presents the 3D object detection results on AGC-V2V. As expected, early fusion achieves slightly better performance than late fusion, benefiting from access to fully aggregated features. Among intermediate fusion methods, Cobevt and V2X-ViT outperform others, indicating their stronger capability in feature aggregation. In contrast, V2VNet performs poorly at higher IoU thresholds, suggesting sensitivity to feature misalignment. Overall, performance remains moderate across all methods, likely due to challenges such as time delays and pose estimation errors in cooperative perception. The late fusion baseline exhibits the lowest accuracy, reflecting error accumulation during post-fusion processing.
@@
Late PointPillars[27] 17.7 13.5
Early PointPillars[27] 19.6 14.1
Intermediate V2VNet [1] 18.4 5.7
Cobevt [28] 46.1 41.7
Where2comm [13] 39.3 31.5
V2X-ViT [4] 44.1 36.6
@@
4.2 Benchmark for VUC 3D object detection
Problem Definition and AGC-VUC. This task aims to improve 3D object detection by leveraging the complementary perception capabilities of ground vehicles and UAVs. Unlike traditional V2V cooperation, the UAV provides a global top-down view that enhances occlusion handling and long-range detection. To support this, we construct the AGC-VUC benchmark by selecting cooperative scenarios in AGC-Drive that involve two vehicles and one UAV. Each scenario lasts approximately 10 seconds, sampled at 10Hz, resulting in 100 cooperative sequences.
Quantitative Results. Table 4 presents the 3D detection performance on AGC-VUC after incorporating the UAV into the cooperative system. A new column, denoted as $\Delta_{\mathrm{UAV}}$, is added to quantify the impact of the UAV’s participation on the perception performance. It can be observed that introducing the UAV improves the performance of all cooperative frameworks across all metrics. Notably, V2VNet achieves the largest improvement, with a $\Delta_{\mathrm{UAV}}$ of +11.5. This is because V2VNet has the lowest baseline performance without the UAV, leaving more room for the UAV to enhance its perception capabilities.
@@ 4.2 Benchmark for VUC 3D object detection
Co-Mode Model V2V V2U
mAP@0.5 mAP@0.7 mAP@0.5 mAP@0.7 $\Delta_{\mathrm{UAV}}$
Intermediate V2VNet [1] 30.5 14.6 40.1 27.9 +11.5
Cobevt [28] 42.3 36.9 42.9 37.5 +0.6
Where2comm [13] 42.6 30.7 44.2 32.0 +1.5
V2X-ViT [4] 38.3 28.7 42.6 33.9 +4.8
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported by the National Science and Technology Major Project (2022ZD0117901), the National Natural Science Foundation of China (62206015, 62376024), the Beijing Natural Science Foundation (L257003), and the Young Scientist Program of the National New Energy Vehicle Technology Innovation Center (Xiamen Branch). We thank the anonymous reviewers for insightful discussions.